OpenStudy (anonymous):

a bank customer(or a crazy chick) has $172 in a bank account. She withdraws $85 per month for the next 3 months. She also writes 4 checks for $45.75 each. How much money should she deposit to ensure that her balanceis at least $25 at the end of the 3 months? U c y i called her the crazy chick? =.=

OpenStudy (anonymous):

172-3(85)-4(45.75)= 25 25-172-3(85)-4(45.75)= 25-25 147-3(85)-4(45.75) 147-255-183. = -$291